# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v2

@@ Abstract
Language models (LMs) risk inadvertently memorizing and divulging sensitive or personally identifiable information (PII) seen in training data, causing privacy concerns. Current approaches to address this issue involve costly dataset scrubbing, or model filtering through unlearning and model editing, which can be bypassed through extraction attacks. We propose REVS, a novel non-gradient-based method for unlearning sensitive information from LMs. REVS identifies and modifies a small subset of neurons relevant for constituent tokens which form sensitive information. To adequately evaluate our method on truly sensitive information, we curate two datasets: an email dataset naturally memorized by Llama-3-8B and GPT-J-6B, and a synthetic social security number dataset that we tune the models to memorize. Compared to other methods, REVS demonstrates superior performance in unlearning sensitive information and robustness to extraction attacks, while retaining underlying model integrity.
@@ 1 Introduction
We aim to prevent models from generating specific token sequences rather than erasing broader conceptual knowledge evaluated in datasets such as WMDP (Li et al., 2024) and TOFU (Maini et al., 2024). To this end, we curate two benchmark datasets containing sensitive information: Emails, containing real private email addresses inadvertently memorized by the model during pretraining, and SSN, a synthetic dataset where we instilled social security numbers into LMs through finetuning. Unlike prior work (Patil et al., 2024), which evaluated editing methods on non-sensitive data, our benchmarks contain actual private information, enabling rigorous assessment of unlearning efficacy and robustness against extraction attacks. We conduct experiments on Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021), which have naturally memorized private email addresses from their training data (The Pile; Gao et al., 2020).
Our experiments demonstrate that REVS outperforms six strong baselines in unlearning sensitive information while maintaining model performance and robustness against extraction attacks. Our main contributions are:
  • We introduce REVS, a novel method for unlearning sensitive information from LM parameters.
  • We curate Emails and SSN, two datasets containing sensitive information.
@@ 2 Problem Setup and Preliminaries
Consider the case where a LM is prompted with text from its training dataset that contains sensitive information, e.g., “Contact David Lewis at ”. If the model has memorized the completion of this prompt, it will generate the relevant email address, such as “lewis.david@email.com”, illustrated in Figure 1. This scenario highlights a prevalent issue with LMs: their ability to memorize can lead to unintended divulgence of sensitive data, in particular personally identifiable information (PII).
Formally, let $\mathcal{D}$ be a dataset, and let $M_{\mathcal{D},\Theta}$ be a LM with parameters $\Theta$ trained on $\mathcal{D}$. We consider the model $M_{\mathcal{D},\Theta}$ to be a privacy risk when the model reproduces a sequence of tokens containing PII, $S$, from $\mathcal{D}$ when prompted with a sequence of tokens $P$. Dataset scrubbing approaches tackle this issue from the data perspective, preventing sensitive information from being stored in the model during training. Machine unlearning approaches take on a post-hoc approach, modifying either all, or just a subset of parameters that contain sensitive information. The goal of these approaches is to efficiently and effectively prevent the model from generating each sequence containing sensitive information $S$ in a way robust to extraction attacks, while at the same time preserving general capabilities of the model (Liu et al., 2024).
@@
We briefly describe the main components in Transformer LMs; more details are found elsewhere (Vaswani et al., 2017; Black et al., 2021). Omitting some technical details, a Transformer LM is made of a sequence of blocks. Each block contains self-attention and multi-layer perceptrons (MLPs) that write and read to the residual stream (Elhage et al., 2021). The MLP is defined as $\text{MLP}(x)$ $=\text{FF}_{2}\sigma(\text{FF}_{1}x)$, and is added to the residual stream. The final layer hidden state is projected to the vocabulary space via the unembedding matrix $U$, followed by a softmax to obtain next token probabilities.
@@ 3 Methodology
To unlearn sensitive information, REVS aims to decrease $r_{t}$ to a desired rank $r_{d}$ by identifying and modifying $FF_{2}$ columns that contribute most to generating the target token $t$. This unlearning process operates in two main phases: localization (Section 3.1), where we identify the layers and neurons ($FF_{2}$ columns) which contain information relevant for generating $t$, and editing (Section 3.2), where we overwrite information relevant for $t$ while not affecting knowledge used for generating other tokens using the pseudoinverse of the unembedding matrix $U^{\dagger}$. We provide a high-level overview of REVS in Algorithm 1.
@@
Algorithm 1 REVS: Rank-based Editing for Vocabulary Suppression
@@ 3.1 Neuron Localization
To effectively unlearn sensitive information while minimizing model disruption, we identify specific model components that strongly contribute to generating token $t$. While information in transformer models is distributed across multiple layers and parameters within them, prior work shows it can be effectively localized (Wu et al., 2023; Dai et al., 2022). Our localization process operates in two steps: identifying relevant layers and selecting a small subset of neurons ($FF_{2}$ columns) within them which contain relevant information.
@@ 3.2 Neuron Editing
    Project back to neuron space: $\vec{n}_{j}^{*}=U^{\dagger}\vec{v}$.
@@ 4.1 Models
Throughout our experiments, we use Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021) – models which we find to have memorized sensitive email addresses, originating from their training on The Pile (Gao et al., 2020).
@@ 4.2 Datasets
Emails consists of examples from The Pile (Gao et al., 2020) containing email addresses naturally memorized by each model during training. We identify $96$ and $288$ sentences for Llama-3-8B and GPT-J-6B, respectively.
Synthetic SSN dataset consists of $200$ sentences containing social security numbers, generated by prompting Claude 3 Sonnet (Anthropic, 2024). Each of the base models was fine-tuned to memorize all instances.
We detail dataset curation in Table 10.
@@ 4.4.2 Model Integrity
where $P_{i}$ is a prompt and $S_{i}$ is the original memorized token sequence, such that prior unlearning $M_{\theta}(P_{i})=S_{i}$.
@@
We evaluate the models performance on the MMLU benchmark (Hendrycks et al., 2021) before and after unlearning to measure degradation in general knowledge capabilities (Gao et al., 2024).
@@ 4.4.3 Extraction Resistance
The underlying intuition is that an attack is successful if the attacker can extract the target in any layer. Notably, compared to previous work on unlearning sensitive information (Patil et al., 2024), our extraction attacks are more comprehensive and stricter by considering more candidate tokens across all model layers.
@@
Logit-Lens Attack (LLA): Considers the top-$k$ and bottom-$k$ tokens in the vocabulary logit vector $\vec{v}$ as candidates, obtained by projecting each layer’s residual hidden state to the vocabulary logits.
@@
Delta Attack (DA): Considers the top-$k$ tokens with the largest absolute changes in the vocabulary logit vector $\vec{v}$ between consecutive layers as candidates.
@@
Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ General.@100 $\uparrow$ Specificity $\uparrow$ MMLU $\uparrow$
SSN Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $100{\pm 0.00}$ 61.0561.0561.05{}61.05
FT-L $\underline{36.98{\pm 11.97}}$ $\underline{63.88{\pm 9.88}}$ $\underline{50.35{\pm 10.76}}$ $24.33{\pm 9.78}$ 60.9960.9960.99{}60.99
MEMIT $24.72{\pm 7.21}$ $30.70{\pm 9.67}$ $23.90{\pm 7.61}$ $22.67{\pm 6.50}$ 61.0261.0261.02{}61.02
Max-Entropy $5.12{\pm 2.13}$ $5.17{\pm 3.00}$ $3.92{\pm 2.33}$ $1.40{\pm 0.60}$ 61.0661.0661.06{}61.06
Head-Projection $2.98{\pm 0.79}$ $3.08{\pm 1.23}$ $2.95{\pm 0.68}$ $4.17{\pm 2.41}$ 61.0661.0661.06{}61.06
RMU $16.42{\pm 9.10}$ $13.47{\pm 8.42}$ $16.67{\pm 10.41}$ $\underline{38.67{\pm 14.92}}$ 60.8360.8360.83{}60.83
NPO-KL $11.95{\pm 4.87}$ $38.78{\pm 18.34}$ $36.13{\pm 16.59}$ $6.33{\pm 3.68}$ 61.0161.0161.01{}61.01
REVS (ours) $\bm{89.58{\pm 1.99}}$ $\bm{98.88{\pm 1.28}}$ $\bm{89.67{\pm 3.78}}$ $\bm{82.17{\pm 5.08}}$ 60.8760.8760.87{}60.87
Emails Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$- $100{\pm 0.00}$ 62.1762.1762.17{}62.17
FT-L $\underline{58.20{\pm 2.42}}$ $55.07{\pm 2.40}$- $\underline{61.91{\pm 4.19}}$ 62.1562.1562.15{}62.15
MEMIT $36.98{\pm 5.78}$ $51.12{\pm 4.85}$- $29.25{\pm 5.97}$ 62.2262.2262.22{}62.22
Max-Entropy $31.45{\pm 5.60}$ $\bm{75.15{\pm 2.44}}$- $20.07{\pm 4.32}$ 62.1162.1162.11{}62.11
Head-Projection $34.05{\pm 5.41}$ $67.05{\pm 2.85}$- $23.13{\pm 4.95}$ 62.1062.1062.10{}62.10
RMU $15.65{\pm 7.62}$ $13.80{\pm 10.05}$- $37.76{\pm 17.07}$ 62.1062.1062.10{}62.10
NPO-KL $33.32{\pm 6.96}$ $22.90{\pm 5.47}$- $\bm{62.59{\pm 9.25}}$ 62.0562.0562.05{}62.05
REVS (ours) $\bm{66.67{\pm 2.42}}$ $\underline{74.20{\pm 2.33}}$- $60.88{\pm 4.92}$ 61.7761.7761.77{}61.77
Table 1: Unlearning Effectiveness and Model Integrity results on Llama-3-8B for $k=100$. Best results bold, second best underlined. REVS is superior on all aggregate metrics, and on almost all individual.
@@ 4.5 Implementation Details
In each experiment, we first unlearn all sensitive sequences from each model and then evaluate the edited model. The Unlearning Score is the harmonic mean of Efficacy, Specificity, and Generality (for the SSN dataset), while the Resistance Score is the harmonic mean of all the extraction attacks. We optimize hyperparameters for each method and dataset, to maximize the Unlearning Score with $k=100$ using a fixed seed (00). We perform cross-validation over six random $/50$ 50 splits (using seeds 11116666). The first half of the dataset was selected as the forget set and evaluated for efficacy, generalization (for the SSN dataset), and extraction resistance, while the second half was used as the retain set to evaluate specificity.
@@ 5 Results
Figure 3: Results for Extraction Resistance on Llama-3-8B for $k=100$. REVS is more robust to extraction attacks. We report the full results in tabular format in Table 2.
Figure 4: Llama-3-8B; Top row: SSN dataset; Bottom row: Email dataset. From left to right: (a+d) efficacy vs candidates size, (b+e) efficacy vs. specificity trade-off, (c+f) Unlearning Score vs. number of targets for Emails (c) and number of targets for SSN (f). Mean results with confidence intervals.
REVS outperforms all baselines on the Unlearning Score and the Resistance Score for the Llama-3-8B model across both datasets (Tables 1 and 2). Across individual metrics, REVS consistently achieves the best or close second-best performance. Results on GPT-J-6B, which follow the same patterns, are in Section A.2.
Efficacy.
On the SSN dataset, REVS significantly outperforms all other baselines in unlearning effectiveness. On the Email dataset, REVS performs is a close second but performs best in terms of the aggregate unlearning score.
Generalization.
When applied to other contexts tuned to generate the same SSN, REVS strongly outperforms alternatives, indicating its ability to comprehensively unlearn generating the target tokens across different contexts.
Specificity.
REVS outperforms other baselines on the SSN dataset. For the Email dataset, REVS is close to the best performers while maintaining best unlearning unlearning score.
General capabilities.
MMLU scores remain consistent across all methods compared to the unedited baseline, indicating that the unlearning process does not significantly impact the model’s core knowledge. Consequently, we emphasize specificity as an important metric that captures delicate local perturbations that global metrics like MMLU might overlook.
@@
As shown in Figures 3, LABEL: and 2, REVS performs best on the Resistance Score in both datasets. On SSN, it is also the best on each metric, while on Emails, REVS is best or within standard deviation in every metric.
@@
Samples of generated text for targeted private email addresses and public email addresses are in Section B.4.

### v3

@@ Abstract
Language models (LMs) risk inadvertently memorizing and divulging sensitive or personally identifiable information (PII) seen in training data, causing privacy concerns. Current approaches to address this issue involve costly dataset scrubbing, or model filtering through unlearning and model editing, which can be bypassed through extraction attacks. We propose REVS, a novel non-gradient-based method for unlearning sensitive information from LMs. REVS identifies and modifies a small subset of neurons relevant for constituent tokens that form sensitive information. To adequately evaluate our method on truly sensitive information, we curate three datasets: email and URL datasets naturally memorized by the models, and a synthetic social security number dataset that we tune the models to memorize. Compared to other methods, REVS demonstrates superior performance in unlearning sensitive information and robustness to extraction attacks, while retaining underlying model integrity.
@@ 1 Introduction
We aim to prevent models from generating specific token sequences rather than erasing broader conceptual knowledge evaluated in datasets such as WMDP (Li et al., 2024) and TOFU (Maini et al., 2024). To this end, we curate three benchmark datasets containing sensitive information: Emails and URLs, containing real addresses inadvertently memorized by the model during pretraining, and SSN, a synthetic dataset where we instilled social security numbers into LMs through finetuning. Unlike prior work (Patil et al., 2024), which evaluated editing methods on non-sensitive data, our benchmarks contain actual private information, enabling rigorous assessment of unlearning efficacy and robustness against extraction attacks. We conduct experiments on Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021), which we find to naturally memorize sensitive information from their training data (The Pile; Gao et al., 2020). Our experiments demonstrate that REVS outperforms six strong baselines in unlearning sensitive information while maintaining model performance and robustness against extraction attacks. Our main contributions are:
  • We introduce REVS, a novel method for unlearning sensitive information from LM.
  • We curate three datasets containing sensitive information, two of which are naturally memorized.
@@ 2 Problem Setup and Preliminaries
Consider the case where a LM is prompted with text from its training dataset that contains sensitive information, e.g., “Contact David Lewis at ”. If the model has memorized the completion of this prompt, it will generate the relevant email address, such as “lewis.david@email.com”, illustrated in Figure 1. This scenario highlights a prevalent issue with LMs: their ability to memorize can lead to unintended divulgence of sensitive data, in particular PII. To address this issue, dataset scrubbing approaches tackle it from a data perspective, preventing sensitive information from being stored in the model during training. Machine unlearning approaches take on a post-hoc approach, modifying either all, or just a subset of parameters that contain sensitive information. The goal of these approaches is to efficiently and effectively prevent the model from generating each sequence containing sensitive information in a way robust to extraction attacks, while preserving general capabilities of the model (Liu et al., 2024).
@@
We briefly describe the main components in Transformer LMs; more details are found elsewhere (Vaswani et al., 2017; Black et al., 2021). Omitting some technical details, a Transformer LM is made of a sequence of blocks. Each block contains self-attention and multi-layer perceptrons (MLPs) that write and read to the residual stream (Elhage et al., 2021). The MLP is defined as $\text{MLP}(\vec{x})$ $=\text{FF}_{2}\sigma(\text{FF}_{1}\vec{x})$, and is added to the residual stream. The final layer hidden state is projected to the vocabulary space via the unembedding matrix $U$, followed by a softmax to obtain next token probabilities.
@@ 3 Methodology
To unlearn sensitive information, REVS aims to decrease $r_{t}$ to a desired rank $r_{d}$ by identifying and modifying $FF_{2}$ columns that contribute most to generating the target token $t$. This unlearning process operates in two main phases: localization (Section 3.1), where we identify the layers and neurons ($FF_{2}$ columns) that contain information relevant for generating $t$, and editing (Section 3.2), where we overwrite information relevant for $t$ while not affecting knowledge used for generating other tokens. We provide a high-level overview of REVS in Algorithm 1.
@@
Algorithm 1 REVS: Rank Editing in the Vocabulary Space
@@ 3.1 Neuron Localization
To effectively unlearn sensitive information while minimizing model disruption, we identify specific model components that strongly contribute to generating token $t$. While information in transformer models is distributed across multiple layers and parameters within them, prior work shows it can be effectively localized (Wu et al., 2023; Dai et al., 2022). Our localization process operates in two steps: identifying relevant layers, and selecting a small subset of neurons ($FF_{2}$ columns) within them that contain relevant information.
@@ 3.2 Neuron Editing
    Project back to neuron space: $\vec{n}_{j}^{*}=U^{\dagger}\vec{v}$, where $U^{\dagger}$ is the pseudoinverse of the unembedding matrix.
@@ 4.1 Models
We use Llama-3-8B (Dubey et al., 2024) and GPT-J-6B (Wang and Komatsuzaki, 2021), which we find have memorized sensitive information from their training on The Pile (Gao et al., 2020).
@@ 4.2 Datasets
Organically memorized data: Identifying naturally memorized sensitive information requires two computationally intensive steps: detecting genuine sensitive information and verifying its memorization. Due to computational constraints, we use a pre-filtered subset of The Pile (Gao et al., 2020) containing 15,000 sentences.111github.com/google-research/lm-extraction-benchmark From this, we identified memorized instances in Llama-3-8B (205 email addresses, 203 URLs) and GPT-J-6B (288 email addresses). Synthetic SSN dataset: We generated 200 sentences containing social security numbers using Claude 3 Sonnet (Anthropic, 2024) and fine-tuned the base models to memorize them. See Appendix F for more details.
@@ 4.4.2 Model Integrity
where $P_{i}$ is a prompt and $S_{i}$ is the original memorized token sequence, such that before unlearning, $M_{\theta}(P_{i})=S_{i}$.
@@
We evaluate the models performance on the MMLU (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) benchmarks before and after unlearning to measure degradation in general knowledge capabilities using lm-eval-harness (Gao et al., 2024).
@@ 4.4.3 Extraction Resistance
An attack is successful if the target is extracted from any layer. Compared to Patil et al. (2024), the extraction attacks used in this work are stricter and more comprehensive, considering more candidate tokens across all layers.
@@
Logit-Lens Attack (LLA): Considers the top-$k$ and bottom-$k$ tokens in the vocabulary logit vector $\vec{v}$ as candidates, obtained by projecting each layer’s residual hidden state to the vocabulary logits:
@@
Delta Attack (DA): Considers the top-$k$ tokens with the largest absolute changes in the vocabulary logit vector $\vec{v}$ between consecutive layers as candidates:
@@
Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ General.@100 $\uparrow$ Specificity $\uparrow$ MMLU $\uparrow$ GSM8K $\uparrow$
SSN Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $100{\pm 0.00}$ 61.0561.0561.05{}61.05 $47.83$
FT-L $\underline{36.98{\pm 11.97}}$ $\underline{63.88{\pm 9.88}}$ $\underline{50.35{\pm 10.76}}$ $24.33{\pm 9.78}$ 60.9960.9960.99{}60.99 $46.62$
MEMIT $24.72{\pm 7.21}$ $30.70{\pm 9.67}$ $23.90{\pm 7.61}$ $22.67{\pm 6.50}$ 61.0261.0261.02{}61.02 $46.17$
Max-Entropy $5.12{\pm 2.13}$ $5.17{\pm 3.00}$ $3.92{\pm 2.33}$ $1.40{\pm 0.60}$ 61.0661.0661.06{}61.06 $47.46$
Head-Projection $2.98{\pm 0.79}$ $3.08{\pm 1.23}$ $2.95{\pm 0.68}$ $4.17{\pm 2.41}$ 61.0661.0661.06{}61.06 $46.92$
RMU $16.42{\pm 9.10}$ $13.47{\pm 8.42}$ $16.67{\pm 10.41}$ $\underline{38.67{\pm 14.92}}$ 60.8360.8360.83{}60.83 $48.21$
NPO-KL $11.95{\pm 4.87}$ $38.78{\pm 18.34}$ $36.13{\pm 16.59}$ $6.33{\pm 3.68}$ 61.0161.0161.01{}61.01 $47.23$
REVS (ours) $\bm{89.58^{\dagger}{\pm 1.99}}$ $\bm{98.88{\pm 1.28}}$ $\bm{89.67{\pm 3.78}}$ $\bm{82.17{\pm 5.08}}$ 60.8760.8760.87{}60.87 $44.20$
Emails Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$- $\bm{100{\pm 0.00}}$ 62.1762.1762.17{}62.17 47.9947.9947.99{}47.99
FT-L $\underline{50.30{\pm 3.04}}$ $52.98{\pm 4.23}$- ${49.25{\pm 8.50}}$ 62.1562.1562.15{}62.15 $50.94$
MEMIT $35.43{\pm 4.30}$ $63.63{\pm 3.50}$- $24.84{\pm 4.20}$ 62.2262.2262.22{}62.22 $50.64$
Max-Entropy $31.08{\pm 3.30}$ $\bm{69.75{\pm 6.30}}$- $20.22{\pm 3.10}$ 62.1162.1162.11{}62.11 $50.64$
Head-Projection $30.80{\pm 3.90}$ $\underline{64.33{\pm 4.90}}$- $20.43{\pm 3.40}$ 62.1062.1062.10{}62.10 $50.19$
RMU $17.47{\pm 3.60}$ $15.08{\pm 5.90}$- $32.58{\pm 16.00}$ 62.1062.1062.10{}62.10 $46.39$
NPO-KL $32.75{\pm 2.70}$ $24.27{\pm 3.00}$- $\underline{50.97{\pm 2.00}}$ 62.0562.0562.05{}62.05 $48.67$
REVS (ours) $\bm{62.37^{\dagger}{\pm 2.30}}$ $59.65{\pm 3.95}$- $65.70{\pm 3.79}$ 61.7761.7761.77{}61.77 47.4647.4647.46{}47.46
URL FT-L $\underline{28.03{\pm 3.95}}$ $\underline{59.13{\pm 7.71}}$- ${18.63{\pm 3.52}}$ $62.14$ $50.72$
MEMIT $17.52{\pm 4.10}$ $34.37{\pm 10.80}$- $11.98{\pm 3.00}$ $62.14$ $49.96$
Max-Entropy $12.78{\pm 3.90}$ $32.88{\pm 7.90}$- $8.06{\pm 2.80}$ $62.19$ $49.50$
Head-Projection $11.28{\pm 3.90}$ $26.32{\pm 8.40}$- $7.30{\pm 2.70}$ $62.14$ $49.81$
RMU $13.48{\pm 6.00}$ $12.22{\pm 12.20}$- $\underline{41.83{\pm 15.30}}$ $62.02$ $49.81$
NPO-KL $17.80{\pm 6.60}$ $10.97{\pm 4.40}$- $\bm{50.87{\pm 14.10}}$ $62.13$ $49.88$
REVS (ours) $\bm{44.25^{\dagger}{\pm 5.01}}$ $\bm{78.22{\pm 6.04}}$- $30.94{\pm 4.11}$ $62.31$ 47.7647.7647.76{}47.76
Table 1: Unlearning Effectiveness and Model Integrity on Llama-3-8B for $k=100$. Best results are bold, second-best underlined. REVS achieves superior Unlearning Score values across all datasets compared to baselines, where indicates statistical significance ($p<0.05$) in Wilcoxon signed-rank test.
@@ 4.5 Implementation Details
In each experiment, we unlearn all sensitive sequences before evaluation. The Unlearning Score is the harmonic mean of efficacy, specificity, and generality (for the SSN dataset), while the Resistance Score is the harmonic mean of all extraction attacks. Hyperparameters (HPs) are optimized per method on the SSN and Email datasets to maximize Unlearning Score with $k=100$, using a fixed seed (00). In the URL dataset, we use the same HPs as the Emails dataset, to assess robustness and generalization across naturally memorized PII types. We cross-validate over 6 random splits (seeds 11116666). Each split has a forget set of $50$ targets in Emails and URLs, while in SSN, it comprises half the data. The forget set is used for efficacy, generalization (SSN), and extraction resistance, while the rest is used for specificity.
@@ 5 Results
Figure 3: Results for Extraction Resistance on Llama-3-8B for $k=100$. REVS is more robust to extraction attacks. We report the full results in tabular format in Table 2 (Section A.2).
Figure 4: Llama-3-8B; Top row: SSN dataset; Middle row: Email dataset; Bottom row: URL dataset. From left to right: (a+d+g) efficacy vs. candidates size, (b+e+h) efficacy vs. specificity trade-off, (c+f+i) Unlearning Score vs. number of targets. Mean results with confidence intervals.
As Table 1 and Figure 3 show, REVS outperforms all baselines in Unlearning Score and is on par or better in Resistance Score, for Llama-3-8B across all datasets. Results on GPT-J-6B follow the same trends (Section A.3).
Effectiveness.
REVS achieves the best efficacy and generality on SSN and URLs and strong efficacy on Emails. As shown in Figure 4, perfect efficacy is possible but compromises specificity. Notably, methods surpassing REVS in efficacy on Emails achieve a lower Unlearning Score due to poor specificity.
Model Integrity.
REVS preserves model integrity while achieving strong Unlearning Score. It attains the best specificity for SSN and Emails and remains competitive on URLs. MMLU and GSM8K scores remain stable across methods, except for a GSM8K drop on SSN, likely due to the numerical nature of the targets and REVS’s significantly higher Effectiveness. Overall, we find that unlearning does not substantially impact the model’s core knowledge. However, as global metrics may miss local perturbations, specificity is particularly important for assessing unlearning.
@@
As shown in Figure 3 and Table 2, REVS achieves the highest Resistance Score across all datasets and models, except for Emails on Llama-3-8B, where it ranks a close second. Notably, Tables 2 and 4 (Section A.2) highlight a clear link between higher Effectiveness and improved Resistance Score. While strong Effectiveness often compromises specificity, REVS maintains both, achieving the best or second-best Resistance Score and demonstrating balanced robust unlearning.
@@
Samples of generated text after unlearning are in Section B.3, showing the model generates coherent alternative text instead of memorized private emails.